# Vid2Fluid: 3D Dynamic Fluid Assets from Single-View Videos with Generative Gaussian Splatting

### v1

@@
Abstract.
The generation of 3D content from single-view images has been extensively studied, but 3D dynamic scene generation with physical consistency from videos remains in its early stages. We propose a novel framework leveraging generative 3D Gaussian Splatting (3DGS) models to extract 3D dynamic fluid objects from single-view videos. The fluid geometry represented by 3DGS is initially generated from single-frame images, then denoised, densified, and aligned across frames. We estimate the fluid surface velocity using optical flow and compute the mainstream of the fluid to refine it. The 3D volumetric velocity field is then derived from the enclosed surface. The velocity field is then converted into a divergence-free, grid-based representation, enabling the optimization of simulation parameters through its differentiability across frames. This process results in simulation-ready fluid assets with physical dynamics closely matching those observed in the source video. Our approach is applicable to various fluid types, including gas, liquid, and viscous fluids, and allows users to edit the output geometry or extend movement durations seamlessly. Our automatic method for creating 3D dynamic fluid assets from single-view videos, easily obtainable from the internet, shows great potential for generating large-scale 3D fluid assets at a low cost.
Figure 1. Vid2Fluid takes a single-view video of fluid as input shown on the left and automatically extract high-quality ready-to-use dynamic fluid asset shown on the right. In case (a), we preserve the camera viewpoint and the appearance of the original fluid while in case (b) the appearance and the viewpoint are modified.
1. Introduction
@@
Despite the growing demand, the research on fluid 3D generation and simulation remains relatively unexplored. To address this need, we have conducted extensive research to understand the requirements and pain points of artists and content creators in the field. Our findings reveal several key insights: (1) Artists desire a tool that can directly generate fluid assets from a given video, preserving the visual and dynamic characteristics of the original content. (2) In most cases, artists are interested in fluid assets with relatively small advancement, which simplifies the problem and reduces the complexity of the simulation. (3) The generated fluid assets should be easily editable and integrable into existing 3D scenes.
Based on these insights, we proposed a novel method, Vid2Fluid, that addresses the unique challenges of extracting 3D dynamic fluid assets from single-view videos. We approximate the fluid with a meaningful geometry represented by 3D Gaussians (3DGS) based on generative methods and ensure the geometry consistency across the frames. The volumetric field is then estimated based on the 3D generated geometry and motion presented in the video. We also design an optimization framework to approximate the simulation parameters of the fluid. Our experiments show that our method produces high-quality results for fluids with normal advancement, making it suitable for a wide range of artistic and practical applications. Moreover, our method also demonstrate high versatility through extensive experiments, showcasing its effectiveness in handling various types of fluids, rendering options, and editable features.
Our main contributions can be summarized as follows:
  • The problem we aim to solve is a novel task with practical real-world applications and user-editable features. We will open-source the implementation of our method.
  • We design the Geometry and Motion Reconstruction module to unify the inconsistent 3DGS generated fluid from different frames of the video into a coherent geometry and accurately estimate the volumetric velocity field of the fluid.
@@
Figure 2. Overview of our framework.
2. Related Work
2.1. 3D Generation with Gaussian Splatting
@@ 2. Related Work
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) adopts a point-based radiance field, using 3D Gaussian primitives to represent scenes. It has emerged as a prevalent research topic in 3D representation (Zhang et al., 2024b; Huang et al., 2024; Yu et al., 2024b; Lu et al., 2024; Yu et al., 2024a) due to its ability to depict high-quality geometry and textures in novel view synthesis.
@@
3DGS provides a new perspective not only for real-time scene reconstruction but also for 3D generation. DreamGaussian (Tang et al., 2023) optimizes a 3DGS through score distillation using a pre-trained text-to-image diffusion model. Recent methods (He et al., 2025; Zhang et al., 2024a; Zhou et al., 2024) investigate directly train diffusion models on Gaussian splats for higher efficiency. However, the direct methods may struggle to handle real-world inputs, like fluid objects, since they are usually trained on synthetic 3D datasets. LGM (Tang et al., 2025) transforms the single-view generation problem into a multi-view 3DGS reconstruction task using a pre-trained single-view to multi-view 2D diffusion model. The most recent transformer-based methods (Zou et al., 2024; Xiang et al., 2024) are proposed to achieve faster generation with higher quality. TriplaneGaussian (Zou et al., 2024) creates a point cloud from a single image and uses a hybrid triplane-Gaussian representation to greatly accelerate the generation. TRELLIS (Xiang et al., 2024) employs rectified flow transformers unified structured latent to get high-quality results.
2.2. Physics-based Fluid Simulation and Dynamic Gaussians
In physics-based simulation, fluids are often simulated using Eulerian grids and/or Lagrangian particles (Bridson, 2015). Resolving fluid dynamics requires solving the simplified version of the Euler equation (Bardos and Titi, 2007) or the complete version of the Navier-Stokes equation (Doering and Gibbon, 1995) regarding viscosity. Methods of solving these equations advanced from giving analytic solutions  (Wejchert and Haumann, 1991; Stam and Fiume, 1993) to later numerical discretization schemes  (Foster and Metaxas, 1996; Tome and McKee, 1994). On the Lagrangian side, Smoothed Particle Hydrodynamics (SPH) methods  (Monaghan, 1992; Becker and Teschner, 2007; Solenthaler and Pajarola, 2009; Ihmsen et al., 2014; Bender and Koschier, 2015) could simulate amorphous objects and particle systems for sub-grid movements. Alternatively to solving force-based equations, Position-based Dynamics  (Müller et al., 2007; Macklin et al., 2016; Macklin and Müller, 2013) was introduced to constrain object positions directly and achieve higher simulation efficiency. For simulation stability, Stam  (Stam, 1999) introduces the concept of semi-Lagrangian advection which brings up the idea of the hybrid field. Hybrid methods  (Harlow, 1988; Brackbill and Ruppel, 1986; Jiang et al., 2015; Fang et al., 2020; Jiang et al., 2016; Hu et al., 2018) combine advantages of Lagrangian and Eulerian schemes, representing fluid by particles while computing dynamics on grids.
@@ 2.2. Physics-based Fluid Simulation and Dynamic Gaussians
The neural radiance expression of objects with dynamics has long been studied for NeRF systems, including deformation capture by canonical and displacement fields  (Park et al., 2021a, b) and dynamics 3D synthesis  (Gao et al., 2021; Li et al., 2021a; Liu et al., 2022; Pumarola et al., 2021; Qiao et al., 2023). Similarly, for the later proposed 3D Gaussian Splatting, dynamics properties have been imposed on the explicit radiance representation of Gaussians. PhysGaussian (Xie et al., 2023) integrated MPM framework to enable reconstructed 3D Gaussians with versatile dynamic behaviors, and GaussianSplashing (Feng et al., 2024) combine Position-based Dynamics with advanced rendering methods to represent Gaussian particles with fluid dynamics and appearances. Instead of giving physical parameters to reconstructed Gaussian subjectively, several works have developed the learning procedure of 3DGS dynamics from references. PhysDreamer (Zhang et al., 2024c) optimizes material parameters through a differentiable MPM simulation to reproduce dynamic Gaussians that behave similarly to input videos. BAGS (Peng et al., 2024) learns the weights of imposed Gaussian Ellipsoid Neural Bounds to animate the reconstructed object according to the input videos. In our work, we embed generated Guardians with physical parameters compatible with APIC method, to align the dynamics behavior of simulated Gaussian particles closest to that of the fluid in the video.
2.3. Velocity Extraction from Videos
@@
We limit the discussion to fluid velocity reconstruction from input videos and images, instead of direct velocity data or experiments. To collect pixel information from videos, optical flow could be effectively utilized in neural reconstructions  (Li et al., 2021b; Du et al., 2021) for general flowing scenes that vary for image styles and object appearances. Targeting on specific fluid textures can noticeably improve learning of dynamics and re-simulation accuracy  (Guan et al., 2022; Liu et al., 2023). For single-view image sequences, vision techniques like tomography can be utilized with some physical priors to achieve visual processing that retains geometric and physical consistency with input fluids (Gregson et al., 2014; Zang et al., 2020). Generally, multi-view inputs are required for 3D velocity reconstruction and novel view fluid synthesis relies on at least sparse viewpoints of the fluid for trained neural networks to infer the unseen sides  (Okabe et al., 2015; Eckert et al., 2018, 2019). Thus, previous work seldom focuses on reconstructing the entire velocity field from the single input as a generative process. To inform the extracted velocity field with physical grounded features, DVP (Deng et al., 2023) successfully encodes vortex features to learn the specific eddy dynamics of fluid but is restricted to 2D profile. Other works  (Franz et al., 2021; Chu et al., 2022; Yu et al., 2024c) extend to 3D space with strict physical constraints and volume rendering, achieving high-fidelity reconstruction while limited to certain fluid types. In our work, we use single-view images to reconstruct the 3D velocity field even for novel views, leveraging the idea of image-conditioned 3D generation.
3. Method
3.1. Overview
@@
Vid2Fluid takes a single-view video of fluid as input and outputs a 3D fluid physics asset. We design a two-stage pipeline to automatically extract crucial information, like the fluid’s geometry, appearances, physical properties, and motion, from the input video to form the final fluid physics. Stage 1: Geometry and Motion Reconstruction. We employ image-conditioned 3D generative models to generate 3DGS for the input frames and preprocess the generated 3DGSs to make them unified, denoised, and dense in the volume. With the 3D geometry information, we estimate the volumetric velocity field of the fluid by interpolating the surface velocity extracted from the frame images into the volume (Sec. 3.3). Stage 2: Simulation Parameter Optimization. To estimate the fluid dynamics in the video through physical evolution, we convert the velocity field on points to a grid-based representation to enable the differentiable grid velocity computation and guide the optimization of simulation parameters, including fluid physical properties and boundary conditions (Sec. 3.4).
3.2. Background
3.2.1. Generative 3D Gaussian Splatting
@@
3D Gaussian Splatting is an explicit radiance-based representation of 3D objects, utilizing high-degree features of shape and color for multi-view synthesis. Due to its differentiable volume rendering capabilities, it can be effectively employed in image-to-3D object generation, enabling the alignment of conditioned image textures. This generation procedure, denoted as $G$, can be described as: $G:\{C\}_{w,h}\rightarrow{\{x,\sigma,A,F\}}_{p}$, where $\{C\}_{w,h}$ represents pixels of the conditioned image at position $(w,h)$. $x$, $\sigma$, $A$, and $F$ denote the position, opacity, covariance matrix, and spherical harmonic features of each Gaussian particle $p$, respectively. Ideally, the input image could be recovered through discretized splatting rendering:
(1) $\sum_{i\in\mathcal{N}}\alpha_{i}\boldsymbol{SH}(r|_{w,h};F_{i})\displaystyle\prod_{j=1}^{i-1}(1-\alpha_{j})\rightarrow\{C\}_{w,h},$
where $\alpha_{i}$ is the product of $\sigma_{i}$ and the projected 2D Gaussian density of where the kernel intersects with the ray in direction $r|_{w,h}$ from the specific pixel at $(w,h)$, and $\boldsymbol{SH}$ denotes the color calculated with features $F_{i}$. This is equivalent to computing a depth-and-opacity weighted average color of particles along the ray direction. Consequently, individual Gaussian points possess adequate geometric and chromatic information, facilitating their representation of point-based models, such as fluids.
3.2.2. Affine Particle in Cell Method and Dynamic 3D Gaussians
@@
Affine Particle-in-Cell (APIC) (Jiang et al., 2015) is a numerical technique for simulating particles in continuous media, enhancing stability and precision through additional computation of local affine velocity fields. In a time step $n$, the particle velocity is first transferred to the grid together with its affine part:
(2) $(mv)_{i}^{n}$ $=\sum_{p}w_{ip}m_{p}[v^{n}_{p}+\mathbf{C}_{p}^{n}(\mathbf{x}_{i}-\mathbf{x}_{p}^{n})]$
$v_{i}^{n}$ $=(mv)_{i}^{n}\big{/}\sum_{p}w_{ip}m_{p},$
@@
where $i$ and $p$ denote grid and particle coordinates respectively. $m$ is the mass, $v$ is the velocity, $\omega$ is the weight function, such as quadratic B-Spline kernel, $\mathbf{C}$ is the affine matrix associated with each particle. Assuming the simulated media is incompressible, the velocity change $\frac{\partial v}{\partial t}$ can be computed on the grid by the following equation:
(3) $\frac{\partial v}{\partial t}=$ $-\frac{\nabla p}{\rho}+\nabla\cdot(2\nu\textbf{S}_{ij})+g$
$\textbf{S}_{ij}=$ $\frac{1}{2}[\nabla\emph{v}+(\nabla\emph{v})^{\mathsf{T}}],$
[…]
(4) $\mathbf{v}_{p}^{n+1}$ $=\sum_{i}w_{ip}(\mathbf{v}_{i}^{n}+\frac{\partial v^{n}_{i}}{\partial t}),$
@@
3.3. Reconstruct 3D Gaussians with Dynamics
3.3.1. Preprocess with Generative Gaussian Splatting
Since our input only contains single-view information, to obtain the 3D

### v2

# 3D Dynamic Fluid Assets from Single-View Videos with Generative Gaussian Splatting

@@
Abstract
While the generation of 3D content from single-view images has been extensively studied, the creation of physically consistent 3D dynamic scenes from videos remains in its early stages.
We propose a novel framework leveraging generative 3D Gaussian Splatting (3DGS) models to extract and re-simulate 3D dynamic fluid objects from single-view videos using simulation methods.
The fluid geometry represented by 3DGS is initially generated and optimized from single-view images, then denoised, densified, and aligned across frames.
We estimate the fluid surface velocity using optical flow, propose a mainstream extraction algorithm to refine it.
The 3D volumetric velocity field is then derived from the velocity of the fluid’s enclosed surface.
The velocity field is therewith converted into a divergence-free, grid-based representation, enabling the optimization of simulation parameters through its differentiability across frames. This process outputs simulation-ready fluid assets with physical dynamics closely matching those observed in the source video.
Our approach is applicable to various liquid fluids, including inviscid and viscous types, and allows users to edit the output geometry or extend movement durations seamlessly. This automatic method for creating 3D dynamic fluid assets from single-view videos, easily obtainable from the internet, shows great potential for generating large-scale 3D fluid assets at a low cost.
[…]
volume: 45issue: 2
1 Introduction
@@
Despite the growing demand, the research on fluid 3D generation and simulation remains relatively unexplored. To address this need, we have conducted extensive research to understand the requirements and pain points of artists and content creators in the field. Our findings reveal several key insights: (1) Artists seek a tool that can directly generate fluid assets from a given video, preserving the original geometry and dynamic characteristics while enabling customizable digital rendering. (2) In most cases, artists are interested in fluid assets with relatively small advancement, which simplifies the problem and reduces the complexity of the simulation. (3) The generated fluid assets should be easily editable and integrable into existing 3D scenes.
Based on these insights, we proposed a novel method that addresses the unique challenges of extracting 3D dynamic fluid assets from single-view videos through physics-based simulation. We approximate the fluid with a meaningful geometry represented by 3D Gaussians (3DGS) based on generative methods and ensure the geometry consistency across the frames. The volumetric field is then estimated based on the 3D generated geometry and motion presented in the video. This process leverages optic flow to analyze the pixel motion in the video, corrected by a physics-derived constraint to retain real fluid features. We also design an optimization framework to approximate the simulation parameters of the fluid that best recovers the dynamics from the video. Our experiments show that the proposed method produces high-quality results for fluids with convectional flow motion, making it suitable for a wide range of artistic and practical applications. Comparisons are demonstrated to show the necessity and efficacy of the optimization process, whereas simply reconstructing precise geometry but manually simulating the fluid produces far inferior results. Moreover, our method also demonstrate high versatility through extensive experiments, showcasing its effectiveness in handling various types of fluids, rendering options, and editable features.
It should be noted that our method does not prioritize precise reconstruction of the visual appearance. First, the target scene is an open system with inflow and outflow, which is different from closed systems like deformable solids where per-vertex tracking is feasible. Second, our single-view input relies on generative inpainting to complete the 3D appearance, where the limited-view guidance provides no information about other perspectives. Finally, the amorphous nature of fluids means the exposure of the inner part easily breaks the surface texture which is more reliably inferred from surface physics. Therefore, we focus more on recovering fluid dynamics by re-simulation to reveal its underlying physics. Implicit representations of fluid motion often fail to preserve physical laws such as incompressibility and momentum conservation. These high-degree-of-freedom phenomena are poorly represented by directly optimizing displacement or velocity fields, which also lack temporal scalability. In contrast, our physics-based approach can generate spatially and temporally coherent 3D fluids that closely match the video dynamics and are also extendable. Moreover, it provides digital creators with an editable and interactive model, as the intuitive physical parameters are straightforward to tune.
The main contributions can be summarized as follows:
  • We propose a novel open-source framework to extract 3D fluid assets from single-view videos, tackling an under-explored problem with practical real-world applications and user-editable features.
  • We design customized geometry and motion reconstruction strategies to produce more coherent geometry and physically faithful velocity field.
@@
2 Related Work
2.1 3D Generation with Gaussian Splatting
@@ 2 Related Work
3D Gaussian Splatting (3DGS) [KKLD23] adopts a point-based radiance field, using 3D Gaussian primitives to represent scenes. It has emerged as a prevalent research topic in 3D representation [ZFS24, HYC24, YSG24, LYX24, YCH24] due to its ability to depict high-quality geometry and textures in novel view synthesis.
@@
3DGS provides a new perspective not only for real-time scene reconstruction but also for 3D generation. DreamGaussian [TRZ23] optimizes a 3DGS through score distillation using a pre-trained text-to-image diffusion model. Recent methods [HCP25, ZCY24, ZZL24] investigate directly training diffusion models on Gaussian splats for higher efficiency. However, the direct methods may struggle to handle real-world inputs, like fluid objects, since they are usually trained on synthetic 3D datasets. LGM [TCC25] transforms the single-view generation problem into a multi-view 3DGS reconstruction task using a pre-trained single-view to multi-view 2D diffusion model. The most recent transformer-based methods [ZYG24, XLX24] are proposed to achieve faster generation with higher quality. TriplaneGaussian [ZYG24] creates a point cloud from a single image and uses a hybrid triplane-Gaussian representation to greatly accelerate the generation. TRELLIS [XLX24] employs rectified flow transformers unifying structured latent to get high-quality results.
2.2 Dynamic Gaussian Reconstruction and Physics-based Fluid Simulation
The neural radiance expression of objects with dynamics has long been studied for NeRF systems, including deformation capture by canonical and displacement fields  [PSB21, PSH21] and dynamics 3D synthesis  [GSKH21, LNSW21a, LCM22, PCPMMN21, QGX23]. Similarly, for the later proposed 3D Gaussian Splatting, dynamics properties have been imposed on the explicit radiance representation of Gaussians. Leveraging the high fitting ability of Gaussian particles, many studies have explored dynamic scene reconstruction guided by image-based losses from video [RXM24, WYF24, YPTW24, GXC24, NRS22, DWD24].
In physics-based simulation, fluids are often simulated using Eulerian grids and/or Lagrangian particles [Bri15], with different consideration of viscosity  [BT07, DG95]. Methods of solving these equations have advanced with numerical discretization schemes  [FM96, TM94]. For simulation stability, Stam  [Sta99] introduces the concept of semi-Lagrangian advection which brings up the idea of the hybrid field. Hybrid methods  [Har88, BR86, JSS15, FHNJ20, JST16, HFG18] combine advantages of Lagrangian and Eulerian schemes, representing fluid by particles while computing dynamics on grids.
@@ 2.2 Dynamic Gaussian Reconstruction and Physics-based Fluid Simulation
Several studies have augmented static Gaussian points with physical parameters to make them animatable. PhysGaussian [XZQ23] integrats MPM framework to enable reconstructed 3D Gaussians with versatile dynamic behaviors, and GaussianSplashing [FFS25] combines Position-based Dynamics with advanced rendering methods to represent Gaussian particles with fluid dynamics and appearances. PhysMotion [TJL24] proposes a framework for animating 3DGS generated from a single-view image. The sequential rendering results are then composited with the input background through an inversion and diffusion process to obtain videos. Instead of giving physical parameters to reconstructed Gaussian subjectively, several works have developed the learning procedure of 3DGS dynamics from references. Simplicits [MSP24] uses a neural field of learnable weights for reduced-order simulation. These works mainly apply to unsupervised simulation. Given the guidance, PhysDreamer [ZYW24] optimizes material parameters through a differentiable MPM simulation to reproduce dynamic Gaussians that behave similarly to input videos. BAGS [PTZ24] learns the weights of imposed Gaussian Ellipsoid Neural Bounds to animate the reconstructed object according to the input videos.
In our work, we embed generated Gaussians with physical parameters compatible with APIC method, to align the dynamic behavior of simulated Gaussian particles closest to that of the fluid in the video.
2.3 Velocity Extraction from Videos
@@
We limit the discussion to fluid velocity reconstruction from input videos and images, instead of direct velocity data or experiments. To collect pixel information from videos, optical flow could be effectively utilized in neural reconstructions  [LNSW21b, DZY21] for general flowing scenes that vary for image styles and object appearances. Targeting on specific fluid textures can noticeably improve learning of dynamics and re-simulation accuracy  [GDWY22, LCN23]. Generally, multi-view inputs are required for 3D velocity reconstruction and novel view fluid synthesis relies on at least sparse viewpoints of the fluid for trained neural networks to infer the unseen sides  [ODAO15, EHT18, EUT19]. Thus, previous work seldom focuses on reconstructing the entire velocity field from the single input as a generative process. To inform the extracted velocity field with physical grounded features, DVP [DYWZ23] successfully encodes vortex features to learn the specific eddy dynamics of fluid but is restricted to 2D profile. Other works  [FST21, CLZ22, YZG24] extend to 3D space with strict physical constraints and volume rendering, achieving high-fidelity reconstruction while limited to certain fluid types. Recently, FluidNexus [GYZW25] has employed the novel-view video synthesizer to reconstruct the velocity field that closely match input frames. In our work, we use single-view images to resimulate the 3D velocity field available for arbitrary perspectives, leveraging the idea of image-conditioned 3D generation.
3 Method
3.1 Overview
Our goal is to extract and re-simulate fluid physics from single-view videos, which are easily accessible online. To achieve this, we adopt single-view generative 3D methods for geometry reconstruction, with point-based generation being particularly suitable for capturing dynamic properties. This motivates our use of 3D Gaussian Splatting (3DGS) and the integration of current generative 3DGS models. The resulting point representations are compatible with most particle-based fluid simulations. While the open-system nature of fluids makes direct point optimization challenging, our method employs a widely used grid-to-point transformation that combines Eulerian and Lagrangian perspectives, thereby simplifying dynamic optimization with fixed grid coordinates.
@@ 3.1 Overview
Our method takes a single-view video of fluid as input and outputs a 3D fluid physics asset. We design a two-stage pipeline to automatically extract crucial information, like the fluid’s geometry, appearances, physical properties, and motion, from the input video to form the final fluid physics. Stage 1: Geometry and Motion Reconstruction. We employ image-conditioned 3D generative models to generate 3DGS for the input frames and preprocess the generated 3DGSs to improve quality – making them unified, denoised, and dense throughout the volume. With the 3D geometry information, we estimate the volumetric velocity field of the fluid using velocity-free projection from the surface velocity extracted from the frame images (Sec. 3.3). Stage 2: Simulation Parameter Optimization. To estimate the fluid dynamics in the video through physical evolution, we convert the velocity field on points to a grid-based representation to enable the differentiable grid velocity computation and guide the optimization of simulation parameters, including fluid physical properties and boundary conditions (Sec. 3.4).
Figure 1: Our framework consists of five stages: (1) generating 3DGS representation from input frames and preprocessing the 3D Gaussians, (2) estimating 2D screen-space velocities using optical flow with mainstream correction, (3) combining with depth information to obtain 3D velocities while extracting terrain geometry, (4) optimizing fluid properties through differentiable simulation, and (5) post-processing for final rendering.
3.2 Background
3.2.1 Generative 3D Gaussian Splatting
@@
3D Gaussian Splatting is an explicit radiance-based representation of 3D objects, utilizing high-degree features of shape and color for multi-view synthesis. Due to its differentiable volume rendering capabilities, it can be effectively employed in image-to-3D object generation, enabling the alignment of conditioned image textures. This generation procedure, denoted as $G$, can be described as: $G:\{C\}_{w,h}\rightarrow{\{x,\sigma,A,F\}}_{p}$, where $\{C\}_{w,h}$ represents pixels of the conditioned image at position $(w,h)$. $x$, $\sigma$, $A$, and $F$ denote the position, opacity, covariance matrix, and spherical harmonic features of each Gaussian particle $p$, respectively. Ideally, the input image could be recovered through discretized splatting rendering: $\sum_{i\in\mathcal{N}}\alpha_{i}\boldsymbol{SH}(r|_{w,h};F_{i})\prod_{j=1}^{i-1}(1-\alpha_{j})\rightarrow\{C\}_{w,h},$ where $\alpha_{i}$ is the product of $\sigma_{i}$ and the projected 2D Gaussian density of where the kernel intersects with the ray in direction $r|_{w,h}$ from the specific pixel at $(w,h)$, and $\boldsymbol{SH}$ denotes the color calculated with features $F_{i}$. This is equivalent to computing a depth-and-opacity weighted average color of particles along the ray direction. Consequently, individual Gaussian points possess adequate geometric and chromatic information, facilitating their representation of point-based models, such as fluids.
3.2.2 Affine Particle in Cell Method and Dynamic 3D Gaussians
@@
Affine Particle-in-Cell (APIC) [JSS15] is a numerical technique for simulating particles in continuous media, enhancing stability and precision through additional computation of local affine velocity fields. In a time step $n$, the particle velocity is first transferred to the grid together with its affine part:
$(m\vec{v})_{i}^{n}$ $=\sum_{p}w_{ip}m_{p}[\mathbf{v}^{n}_{p}+\mathbf{C}_{p}^{n}(\mathbf{x}_{i}-\mathbf{x}_{p}^{n})]$ (1)
$\vec{v}_{i}^{n}$ $=(m\vec{v})_{i}^{n}\big/\sum_{p}w_{ip}m_{p},$
@@
where $i$ and $p$ denote grid and particle coordinates respectively. $m$ is the mass, $\vec{v}_{i}$ is the grid velocity and $\mathbf{v}_{p}$ is the particle velocity. $\omega$ is the weight function, such as quadratic B-Spline kernel, and $\mathbf{C}$ is the affine matrix associated with each particle. Assuming the simulated media is incompressible, the velocity change $\frac{\partial\vec{v}}{\partial t}$ can be computed on the grid by the following equation:
$\frac{\partial\vec{v}}{\partial t}=$ $-\frac{\nabla p}{\rho}+\nabla\cdot(2\nu\textbf{S}_{ij})+g$ (2)
$\textbf{S}_{ij}=$ $\frac{1}{2}[\nabla\vec{v}+(\nabla\vec{v})^{\mathsf{T}}],$
where $\nabla p$ is the gradient of pressure, $\rho$ is the density, $\nabla\cdot$ is the divergence operator, $g$ is the acceleration of body forces like gravity. $\nu$ denotes kinematic viscosity, modeling the viscous stress by the strain rate $\textbf{S}_{i,j}$ which is the deviatoric part of the velocity gradient. Here the splitting scheme is often employed to decouple this partial differential equation (PDE) for robust and efficient solve [Cho67]. The body force is initially applied explicitly, followed by projecting the velocity to be divergence-free by solving the Poisson Pressure Equation. Subsequently, the viscous stress is computed from the strain rate. We refer the readers to [Bri15] for more details. The updated grid velocity is then transferred back to the particles as states of the next time step $n+1$, and the corresponding affine matrix is simultaneously updated:
$\mathbf{v}_{p}^{n+1}$ $=\sum_{i}w_{ip}\big(\vec{v}_{i}^{n}+\big(\frac{\partial\vec{v}_{i}}{\partial t}\big)^{n}\Delta t\big),$ (3)
@@
3.3 Reconstruct 3D Gaussians with Dynamics
3.3.1 Preprocess with Generative Gaussian Splatting